# Universal Latent Homeomorphic Manifolds: Cross-Domain Representation Learning via Homeomorphism Verification

### v1

@@ Abstract
We present the Universal Latent Homeomorphic Manifold (ULHM), a framework that unifies semantic representations (e.g., human descriptions, diagnostic labels) and observation-driven machine representations (e.g., pixel intensities, sensor readings) into a single latent structure. Despite originating from fundamentally different pathways, both modalities capture the same underlying reality. We establish homeomorphism, a continuous bijection preserving topological structure, as the mathematical criterion for determining when latent manifolds induced by different semantic-observation pairs can be rigorously unified. This criterion provides theoretical guarantees for three critical applications: (1) semantic-guided sparse recovery from incomplete observations, (2) cross-domain transfer learning with verified structural compatibility, and (3) zero-shot compositional learning via valid transfer from semantic to observation space. Our framework learns continuous manifold-to-manifold transformations through conditional variational inference, avoiding brittle point-to-point mappings. We develop practical verification algorithms, including trust, continuity, and Wasserstein distance metrics, that empirically validate homeomorphic structure from finite samples. Experiments demonstrate: (1) sparse image recovery from 5% of CelebA pixels and MNIST digit reconstruction at multiple sparsity levels, (2) cross-domain classifier transfer achieving 86.73% accuracy from MNIST to Fashion-MNIST without retraining, and (3) zero-shot classification on unseen classes achieving 89.47% on MNIST, 84.70% on Fashion-MNIST, and 78.76% on CIFAR-10. Critically, the homeomorphism criterion correctly rejects incompatible datasets, preventing invalid unification and providing a feasible way to principled decomposition of general foundation models into verified domain-specific components.
@@ 2.2.3 Multi-Objective Training
To learn a shared latent manifold rather than separate manifolds for each modality, we enforce that different modalities from the same sample map to nearby points:
$\mathcal{L}_{\text{consist}}=\sum_{i=1}^{N}\left[\sum_{\begin{subarray}{c}m,m^{\prime}\in\mathcal{M}_{i}^{x}\\ m<m^{\prime}\end{subarray}}\left(1-\frac{z_{i}^{(m)}\cdot z_{i}^{(m^{\prime})}}{\|z_{i}^{(m)}\|_{2}\|z_{i}^{(m^{\prime})}\|_{2}}\right)+\sum_{\begin{subarray}{c}m\in\mathcal{M}_{i}^{x}\\ l\in\mathcal{M}_{i}^{s}\end{subarray}}\left(1-\frac{z_{i}^{(m)}\cdot z_{i}^{(l)}}{\|z_{i}^{(m)}\|_{2}\|z_{i}^{(l)}\|_{2}}\right)\right]$ (11)
where $z_{i}^{(m)}=E_{x}^{(m)}(x_{i}^{(m)};\phi)$ and $z_{i}^{(l)}=E_{s}^{(l)}(s_{i}^{(l)};\phi)$. For image and graph learning reconstruction tasks with sparse observations, we use cosine similarity instead of Euclidean distance to allow directional alignment while permitting magnitude differences between sparse and complete encodings. In contrast, for homogeneous multi-modal scenarios where all modalities provide complete observations with comparable information density (e.g., multi-view imaging, synchronized sensor fusion, or fully-paired text-image data), Euclidean distance is more appropriate as magnitude differences carry semantic meaning. Geometrically, this loss constrains embeddings from all available modalities to overlap significantly in the latent space, creating a unified latent manifold $\mathcal{M}_{z}$.
@@ 2.3.1 Theoretical Analysis
Theorem 1 (Topological Unification).
We denote the data manifold in observation space for dataset $k$ as $\mathcal{M}_{x}^{[k]}$ and the corresponding latent manifold as $\mathcal{M}_{z}^{[k]}$. Let $\{\mathcal{M}_{x}^{[k]}\}_{k=1}^{K}$ be disjoint data manifolds mapped to latent manifolds $\{\mathcal{M}_{z}^{[k]}\}_{k=1}^{K}$ by encoders $\{E^{[k]}\}_{k=1}^{K}$. The union $\mathcal{U}=\bigcup_{k=1}^{K}\mathcal{M}_{z}^{[k]}$ is a valid topological unification preserving the structure of all datasets if:
@@
    Local Preservation: Each $E^{[k]}:\mathcal{M}_{x}^{[k]}\to\mathcal{M}_{z}^{[k]}$ is a homeomorphism.
@@
    Global Connectivity: The intersection graph is connected, and overlaps are non-trivial: $\nu(\mathcal{M}_{z}^{[i]}\cap\mathcal{M}_{z}^{[j]})>0$ for adjacent $i,j$, where $\nu(\cdot)$ denotes the Lebesgue measure on $\mathbb{R}^{d}$.
Theorem 2 (Bi-Lipschitz Sufficiency).
If the encoder $E:\mathcal{M}_{x}\to\mathcal{M}_{z}$ is $(c_{1},c_{2})$-bi-Lipschitz with $c_{1}>0$, then $E$ is a homeomorphism onto its image. Thus, constraining the Lipschitz constants satisfies Condition (i) of Theorem 1.
2.3.2 Empirical Verification Protocol
Direct computation of the bi-Lipschitz constants $c_{1},c_{2}$ and overlap measure $\mu$ is intractable. We bridge theory and practice through computable metrics that detect violations of Theorems 2 and 1 across structural, geometric, and homeomorphic dimensions.
Proposition 1 (Metric-Theoretical Bridge).
Violations of the bi-Lipschitz condition (Theorem 2) and the Manifold Unification condition (Theorem 1) manifest in discrete metrics via a hierarchical dependency, conditioned on sampling density:
  1. 1.
    Structural Fragmentation: If $\beta_{0}(\mathcal{Z}_{\text{total}})>1$ while $W_{2}$ is high, the manifold contains disjoint components, indicating a failure to learn cross-domain transitions. A slightly elevated $\beta_{0}$ under low $W_{2}$ typically indicates finite sampling artifacts rather than model failure.
  2. 2.
    Geometric Misalignment: If $\beta_{0}\approx 1$ but the pairwise distributions $\hat{\mathbb{P}}_{z}^{[i]}$ and $\hat{\mathbb{P}}_{z}^{[j]}$ are globally distant (large $W_{2}$), the domains occupy distinct regions of the shared space with insufficient density overlap $\mu$.
@@ 2.3.2 Empirical Verification Protocol
    Manifold Collapse: If unification holds but the encoder fails to be injective ($c_{1}\to 0$), distinct semantic neighborhoods merge, strictly reducing the Trust Score $\tau_{t}$.
@@
    Inconsistent Mapping: If the encoder is injective but semantic-observation correspondences are violated, Alignment Error $\tau_{a}$ increases.
@@
    Structural Incoherence: If the local neighborhood variance is high relative to the mean distance, the manifold lacks smoothness. This is detected by a drop in Continuity $\tau_{c}$.
@@ 3.3 Zero-Shot Learning via Manifold Learning
Nearest-centroid classification succeeds in the zero-shot regime because the homeomorphic encoder maps the entire data manifold $\mathcal{M}_{\mathcal{X}}$ (including unseen classes) to the latent manifold $\mathcal{M}_{z}$ while preserving topological invariants. The classifier $C_{\theta}$ trained only on seen classes learns decision boundaries in a subspace of $\mathcal{M}_{z}$, but the homeomorphic structure ensures that geometric relationships extend globally. Unseen class samples naturally cluster at their correct centroids through preserved topology, enabling classification without supervision. If topology preservation fails (e.g., manifold collapse where $c_{1}\to 0$, or tearing where continuity is violated), the Trust metric (Proposition 1) drops below threshold, indicating that class boundaries become unreliable. In such cases, zero-shot classification produces arbitrary predictions, as unseen class samples no longer cluster coherently in latent space.
@@ Appendix A Mathematical Proofs
A.1 Mathematical Preliminaries
To establish the continuity of the global inverse mapping across unified latent supports, we rely on the following fundamental result from general topology.
Lemma 1 (Pasting Lemma for Open Sets). Let $X$ be a topological space such that $X=\bigcup_{\alpha\in A}U_{\alpha}$, where each $U_{\alpha}$ is an open set in $X$. Let $Y$ be another topological space. If for each $\alpha\in A$, there exists a continuous map $f_{\alpha}:U_{\alpha}\to Y$ such that for every pair $\alpha,\beta\in A$, the functions agree on their intersection $f_{\alpha}|_{U_{\alpha}\cap U_{\beta}}=f_{\beta}|_{U_{\alpha}\cap U_{\beta}}$, then there exists a unique continuous map $f:X\to Y$ such that $f|_{U_{\alpha}}=f_{\alpha}$ for all $\alpha\in A$.
A.2 Proof of Theorem 1 (Topological Unification)
Theorem 1. Let $\{\mathcal{M}_{x}^{[k]}\}_{k=1}^{K}$ be disjoint data manifolds mapped to latent manifolds $\{\mathcal{M}_{z}^{[k]}\}_{k=1}^{K}$ by encoders $\{E^{[k]}\}_{k=1}^{K}$. The union $\mathcal{U}=\bigcup_{k=1}^{K}\mathcal{M}_{z}^{[k]}$ is a valid topological unification preserving the structure of all datasets if: (i) Each $E^{[k]}$ is a homeomorphism (Local Preservation), and (ii) The intersection graph is connected with non-trivial overlaps $\nu(\mathcal{M}_{z}^{[i]}\cap\mathcal{M}_{z}^{[j]})>0$ (Global Connectivity).
@@ A.2 Proof of Theorem 1 (Topological Unification)
We establish the topological integrity of the unified latent space by constructing a global inverse map $G:\mathcal{U}\to\bigsqcup_{k=1}^{K}\mathcal{M}_{x}^{[k]}$ and invoking the Pasting Lemma for Open Sets.
  1. 1.
    Setup: Let $\mathcal{U}$ be equipped with the subspace topology inherited from the Euclidean distance on $\mathbb{R}^{d}$. We define the global inverse map $G:\mathcal{U}\to\bigcup_{k=1}^{K}\mathcal{M}_{x}^{[k]}$ piecewise by setting $G(z)=(E^{[k]})^{-1}(z)$ for any $z\in\mathcal{M}_{z}^{[k]}$.
  2. 2.
    Well-Definedness on Overlaps: Consider a point $z$ in the overlap region $z\in\mathcal{M}_{z}^{[i]}\cap\mathcal{M}_{z}^{[j]}$. Condition (ii) implies that the geometric realization of the overlap is consistent. Specifically, since $E^{[i]}$ and $E^{[j]}$ represent the encoding of the identical underlying semantic concept in the overlap region, the transition map $\psi_{ij}=E^{[j]}\circ(E^{[i]})^{-1}$ correctly identifies the corresponding points. Thus, $G(z)$ consistently maps to the unique preimage in the disjoint union of domain manifolds, ensuring the global mapping is well-defined across the unified support.
  3. 3.
    Continuity (Local Preservation): Condition (ii) ($\nu>0$) implies that the overlap regions contain open sets. Consequently, the collection $\{\mathcal{M}_{z}^{[k]}\}_{k=1}^{K}$ forms an open cover of the unified space $\mathcal{U}$. Condition (i) states that each individual restriction $E^{[k]}:\mathcal{M}_{x}^{[k]}\to\mathcal{M}_{z}^{[k]}$ is a homeomorphism, which directly implies that each local inverse $(E^{[k]})^{-1}$ is continuous. By Lemma 1 (Pasting Lemma), since each local restriction of the global inverse map is continuous ($G|_{\mathcal{M}_{z}^{[k]}}=(E^{[k]})^{-1}$), the entire global inverse map $G$ is continuous on $\mathcal{U}$.
  4. 4.
    Bijectivity and Global Homeomorphism: The forward map $E$ is surjective onto $\mathcal{U}$ by construction. Injectivity is guaranteed by the injectivity of the local homeomorphisms $E^{[k]}$ and the disjoint nature of the input domain manifolds $\mathcal{M}_{x}^{[k]}$.
Since the global map is a bijection, and both the forward mapping (composed of continuous local pieces) and the global inverse map $G$ are continuous, the unification $\mathcal{U}$ is homeomorphic to the input data structure. This ensures that zero-shot attribute composition and cross-modal transfer remain valid within the unified manifold. ∎
A.3 Proof of Theorem 2 (Bi-Lipschitz Sufficiency)
Theorem 2. If the encoder $E:\mathcal{M}_{x}\to\mathcal{M}_{z}$ is $(c_{1},c_{2})$-bi-Lipschitz with $c_{1}>0$, then $E$ is a homeomorphism onto its image. Thus, constraining the Lipschitz constants satisfies Condition (i) of Theorem 1.

### v2

# Universal Latent Homeomorphic Manifolds: A Framework for Cross-Domain Representation Unification

@@ Abstract
We present the Universal Latent Homeomorphic Manifold (ULHM), a framework that unifies semantic representations (e.g., human descriptions, diagnostic labels) and observation-driven machine representations (e.g., pixel intensities, sensor readings) into a single latent structure. Despite originating from fundamentally different pathways, both modalities capture the same underlying reality. We establish homeomorphism, a continuous bijection preserving topological structure, as the mathematical criterion for determining when latent manifolds induced by different semantic-observation pairs can be rigorously unified. This criterion provides theoretical guarantees for three critical applications: (1) semantic-guided sparse recovery from incomplete observations, (2) cross-domain transfer learning with verified structural compatibility, and (3) zero-shot compositional learning via valid transfer from semantic to observation space. Our framework learns continuous manifold-to-manifold transformations through conditional variational inference, avoiding brittle point-to-point mappings. We develop practical verification algorithms, including trust, continuity, and Wasserstein distance metrics, that empirically validate homeomorphic structure from finite samples. Experiments demonstrate: (1) sparse image recovery from 5% of CelebA pixels and MNIST digit reconstruction at multiple sparsity levels, (2) cross-domain classifier transfer achieving 86.73% accuracy from MNIST to Fashion-MNIST without retraining, and (3) zero-shot classification on unseen classes achieving 89.47% on MNIST, 84.70% on Fashion-MNIST, and 78.76% on CIFAR-10. Critically, the homeomorphism criterion determines when different semantic-observation pairs share compatible latent structure, enabling principled unification into universal representations and providing a mathematical foundation for decomposing general foundation models into domain-specific components.
@@ 2.2.3 Multi-Objective Training
To learn a shared latent manifold rather than separate manifolds for each modality, we enforce that different modalities from the same sample map to nearby points through a composite consistency objective combining both angular and magnitude alignment:
$\mathcal{L}_{\text{consist}}=$ $\lambda_{\text{cos}}\sum_{i=1}^{N}\Bigg[\sum_{\begin{subarray}{c}m,m^{\prime}\in\mathcal{M}_{i}^{x}\\ m<m^{\prime}\end{subarray}}\left(1-\frac{z_{i}^{(m)}\cdot z_{i}^{(m^{\prime})}}{\|z_{i}^{(m)}\|_{2}\|z_{i}^{(m^{\prime})}\|_{2}}\right)+\sum_{\begin{subarray}{c}m\in\mathcal{M}_{i}^{x}\\ l\in\mathcal{M}_{i}^{s}\end{subarray}}\left(1-\frac{z_{i}^{(m)}\cdot z_{i}^{(l)}}{\|z_{i}^{(m)}\|_{2}\|z_{i}^{(l)}\|_{2}}\right)\Bigg]$
$+\lambda_{\text{eucl}}\sum_{i=1}^{N}\Bigg[\sum_{\begin{subarray}{c}m,m^{\prime}\in\mathcal{M}_{i}^{x}\\ m<m^{\prime}\end{subarray}}\|z_{i}^{(m)}-z_{i}^{(m^{\prime})}\|_{2}^{2}+\sum_{\begin{subarray}{c}m\in\mathcal{M}_{i}^{x}\\ l\in\mathcal{M}_{i}^{s}\end{subarray}}\|z_{i}^{(m)}-z_{i}^{(l)}\|_{2}^{2}\Bigg]$ (11)
where $z_{i}^{(m)}=E_{x}^{(m)}(x_{i}^{(m)};\phi)$ and $z_{i}^{(l)}=E_{s}^{(l)}(s_{i}^{(l)};\phi)$, and $\lambda_{\text{cos}},\lambda_{\text{eucl}}\geq 0$ control the strength of angular and magnitude-based alignment, respectively. The cosine similarity term (weighted by $\lambda_{\text{cos}}$) enables directional alignment while permitting magnitude differences, which is essential for reconstruction tasks with sparse observations where information density varies significantly between modalities. The Euclidean distance term (weighted by $\lambda_{\text{eucl}}$) provides additional geometric constraints on embedding magnitudes, ensuring tighter clustering in scenarios where magnitude differences carry semantic meaning. For homogeneous multi-modal scenarios with comparable information density (e.g., multi-view imaging, synchronized sensor fusion, or fully-paired text-image data), we set both $\lambda_{\text{cos}},\lambda_{\text{eucl}}>0$ to enforce both angular and magnitude alignment. For heterogeneous scenarios with asymmetric information content (e.g., sparse vs. complete observations), we set $\lambda_{\text{eucl}}=0$ and $\lambda_{\text{cos}}>0$ to prioritize directional consistency. Geometrically, this composite loss constrains embeddings from all available modalities to overlap significantly in the latent space, creating a unified latent manifold $\mathcal{M}_{z}$.
@@ 2.3.1 Theoretical Analysis
Theorem 1.
(Topological Unification via Latent Identifications). Let $\{\mathcal{M}_{x}^{[k]}\}_{k=1}^{K}$ be pairwise disjoint topological spaces with continuous encoders $E^{[k]}:\mathcal{M}_{x}^{[k]}\to\mathbb{R}^{d}$. Define the latent images $\mathcal{M}_{z}^{[k]}:=E^{[k]}(\mathcal{M}_{x}^{[k]})$ and unified latent support $\mathcal{U}:=\bigcup_{k=1}^{K}\mathcal{M}_{z}^{[k]}\subset\mathbb{R}^{d}$. Let $X:=\bigsqcup_{k=1}^{K}\mathcal{M}_{x}^{[k]}$ (disjoint union, coproduct topology) and define $E:X\to\mathcal{U}$ by $E|_{\mathcal{M}_{x}^{[k]}}=E^{[k]}$. Define $x\sim x^{\prime}$ iff $E(x)=E(x^{\prime})$. Then the induced map $\widetilde{E}:X/\!\sim\ \to\ \mathcal{U}$ is a homeomorphism provided:
  1. (i)
    (Local Homeomorphism) Each $E^{[k]}$ is a homeomorphism onto $\mathcal{M}_{z}^{[k]}$ (subspace topology from $\mathbb{R}^{d}$).
  2. (ii)
    (Coherence) For all $\subseteq\mathcal{U}: VV$ is open in $\mathcal{U}$ iff $(E^{[k]})^{-1}(V)$ is open in $\mathcal{M}_{x}^{[k]}$ for all $k$.
Theorem 2 (Bi-Lipschitz Sufficiency).
Let $(\mathcal{M}_{x},d_{X})$ and $(\mathcal{M}_{z},d_{Z})$ be metric spaces with induced topologies. If encoder $E:\mathcal{M}_{x}\to\mathcal{M}_{z}$ is $(c_{1},c_{2})$-bi-Lipschitz with $c_{1}>0$, then $E$ is a homeomorphism onto its image, satisfying Condition (i) of Theorem 1.
2.3.2 Empirical Verification Protocol
Direct computation of the bi-Lipschitz constants $c_{1},c_{2}$ and verification of the open cover condition (Theorem 1, Condition ii) is intractable for real datasets. We bridge theory and practice through a hierarchical verification framework that employs computable metrics to detect violations of the theoretical conditions across structural, geometric, and homeomorphic dimensions.
Verification Strategy. While Theorems 1 and 2 provide guarantees under ideal conditions (continuous maps, infinite samples), we verify their empirical satisfaction through complementary metrics organized in a hierarchical dependency structure. Each metric targets a specific failure mode:
@@ 2.3.2 Empirical Verification Protocol
    Structural Fragmentation: $\beta_{0}(\mathcal{Z}_{\text{total}})>1$ with high pairwise $W_{2}$ indicates the latent space contains disjoint components, violating the unified manifold assumption. A slightly elevated $\beta_{0}$ under low $W_{2}$ typically reflects finite-sample artifacts rather than fundamental unification failure.
@@
    Geometric Misalignment: $\beta_{0}\approx 1$ but large $W_{2}(\hat{\mathbb{P}}_{z}^{[i]},\hat{\mathbb{P}}_{z}^{[j]})$ indicates domains occupy distant regions with insufficient density overlap, violating the open cover requirement (Theorem 1, Condition ii).
@@
    Manifold Collapse: Low Trust Score $\tau_{t}$ indicates distinct semantic neighborhoods merge in latent space, signaling violation of the bi-Lipschitz lower bound ($c_{1}\to 0$) and thus failure of injectivity (Theorem 2).
@@
    Structural Incoherence: Low Continuity $\tau_{c}$ indicates the encoder tears or folds the manifold—neighbors in input space become distant in latent space—suggesting violation of smooth homeomorphism.
@@
    Inconsistent Mapping: High Alignment Error $\tau_{a}$ indicates different modalities representing the same entity map to distant latent points, violating cross-modal semantic consistency.
These failure modes form a hierarchical dependency: global unification (items 1–2) is necessary for local preservation (items 3–4) to be meaningful, and both are necessary for semantic coherence (item 5).
@@
2.3.3 Theoretical Connection Between ULHM Objectives and Verification Metrics
The ULHM training objectives (Eq. 8) are deliberately designed to enforce the bi-Lipschitz homeomorphism conditions verified by our empirical metrics. We establish the logical chain connecting each training component to its corresponding verification criterion.
Preventing Manifold Collapse (Trust Score $\tau_{t}$). The reconstruction losses $\mathcal{L}_{\text{recon}}^{x}$ and $\mathcal{L}_{\text{recon}}^{s}$ enforce that the encoder must preserve sufficient information to faithfully reconstruct the original inputs. If the encoder were to collapse distinct semantic classes to the same latent region (violating the bi-Lipschitz lower bound $c_{1}>0$), reconstruction quality would degrade as the decoder cannot disambiguate merged representations. This information-preserving pressure maintains local injectivity, which manifests empirically as high Trust scores—semantic neighbors in data space remain semantic neighbors in latent space without spurious mixing.
Preserving Local Topology (Continuity $\tau_{c}$). The local consistency loss $\mathcal{L}_{\text{local}}$ explicitly minimizes distances between latent codes of $\kappa$-nearest neighbors in data space. This directly implements the continuity requirement of homeomorphism: points close in $\mathcal{M}_{x}$ must map to points close in $\mathcal{M}_{z}$. By construction, this objective optimizes precisely for the Continuity metric, ensuring the encoder does not tear the manifold by mapping nearby data points to distant latent regions.
Achieving Geometric Unification (Wasserstein Distance $W_{2}$ and Alignment Error $\tau_{a}$). The cross-modal consistency loss $\mathcal{L}_{\text{consist}}$ operates at two complementary levels. The cosine similarity component (weighted by $\lambda_{\text{cos}}$) enforces directional alignment, ensuring different modalities representing the same semantic content point toward the same region in latent space, which directly reduces the distributional distance between domains measured by $W_{2}$. The Euclidean distance component (weighted by $\lambda_{\text{eucl}}$) further constrains the magnitude alignment, minimizing $\tau_{a}$ by ensuring paired semantic-observation samples $(s_{i},x_{i})$ map to nearby coordinates. Together, these components enforce the global connectivity condition (Theorem 1) by ensuring non-trivial overlap $\nu(\mathcal{M}_{z}^{[i]}\cap\mathcal{M}_{z}^{[j]})>0$ between domain manifolds.
Ensuring Topological Unity (Betti Number $\beta_{0}$). The combination of reconstruction, consistency, and local preservation losses creates a unified manifold structure. Reconstruction losses prevent the latent space from fragmenting into disconnected task-specific subspaces by requiring all modalities to decode successfully from shared representations. Cross-modal consistency explicitly bridges domains by constraining them to occupy overlapping regions. The perceptual loss $\mathcal{L}_{\text{percep}}$ further reinforces structural coherence by matching high-level feature statistics. These combined pressures ensure the latent space forms a single connected component ($\beta_{0}=1$) rather than disjoint manifolds.
In summary, the ULHM objective function is not an ad hoc combination of losses, but a principled implementation of the bi-Lipschitz homeomorphism criterion: $\mathcal{L}_{\text{recon}}$ enforces the lower bound $c_{1}>0$ (preventing collapse), $\mathcal{L}_{\text{local}}$ enforces the upper bound $c_{2}<\infty$ (preventing tearing), and $\mathcal{L}_{\text{consist}}$ enforces global connectivity (enabling unification). Each verification metric subsequently validates whether these theoretical conditions hold empirically on finite samples.
@@ 3.3 Zero-Shot Learning via Manifold Learning
Nearest-centroid classification succeeds in the zero-shot regime because the homeomorphic encoder maps the entire data manifold $\mathcal{M}_{\mathcal{X}}$ (including unseen classes) to the latent manifold $\mathcal{M}_{z}$ while preserving topological invariants. The classifier $C_{\theta}$ trained only on seen classes learns decision boundaries in a subspace of $\mathcal{M}_{z}$, but the homeomorphic structure ensures that geometric relationships extend globally. Unseen class samples naturally cluster at their correct centroids through preserved topology, enabling classification without supervision. If topology preservation fails (e.g., manifold collapse where $c_{1}\to 0$, or tearing where continuity is violated), the Trust metric (Proposition LABEL:prop:bridge) drops below threshold, indicating that class boundaries become unreliable. In such cases, zero-shot classification produces arbitrary predictions, as unseen class samples no longer cluster coherently in latent space.
@@ Appendix A Mathematical Proofs
A.1 Proof of Theorem 1 (Topological Unification)
Theorem 1 (Topological Unification via Latent Identifications). Let $\{\mathcal{M}_{x}^{[k]}\}_{k=1}^{K}$ be pairwise disjoint topological spaces and let $E^{[k]}:\mathcal{M}_{x}^{[k]}\to\mathbb{R}^{d}$ be continuous maps. Define the latent images $\mathcal{M}_{z}^{[k]}:=E^{[k]}(\mathcal{M}_{x}^{[k]})$ and let
$\mathcal{U}:=\bigcup_{k=1}^{K}\mathcal{M}_{z}^{[k]}\subset\mathbb{R}^{d}$
be equipped with the subspace topology $\tau_{\mathrm{sub}}$ induced from $\mathbb{R}^{d}$. Let $X:=\bigsqcup_{k=1}^{K}\mathcal{M}_{x}^{[k]}$ be the disjoint union equipped with the coproduct topology, and define the global encoder $E:X\to\mathcal{U}$ by $E|_{\mathcal{M}_{x}^{[k]}}=E^{[k]}$. Define an equivalence relation $\sim$ on $X$ by
$x\sim x^{\prime}\quad\Longleftrightarrow\quad E(x)=E(x^{\prime}).$
Assume:
  1. (i)
    (Local Homeomorphism) Each $E^{[k]}$ is a homeomorphism onto its image $\mathcal{M}_{z}^{[k]}$ (with $\mathcal{M}_{z}^{[k]}$ carrying the subspace topology inherited from $\mathbb{R}^{d}$ (equivalently, from $\mathcal{U}$)).
  2. (ii)
    (Coherence / Final-Topology Condition) The subspace topology on $\mathcal{U}$ coincides with the final topology induced by the family $\{E^{[k]}\}_{k=1}^{K}$, i.e., for every $V\subseteq\mathcal{U}$,
    $V\in\tau_{\mathrm{sub}}\quad\Longleftrightarrow\quad(E^{[k]})^{-1}(V)\ \text{is open in}\ \mathcal{M}_{x}^{[k]}\ \text{for all}\ k.$
Then the induced map $\widetilde{E}:X/\!\sim\ \to\ \mathcal{U}$ is a homeomorphism. Hence $\mathcal{U}$ is a valid topological unification of the datasets up to the identifications encoded by overlaps.
@@ A.1 Proof of Theorem 1 (Topological Unification)
We briefly clarify the roles of the two assumptions. Condition (i) ensures that each component $\mathcal{M}_{x}^{[k]}$ is embedded faithfully into latent space, so that no local collapse or topological distortion occurs within a modality and any overlap in latent space corresponds to a genuine identification of points. Condition (ii) guarantees global coherence: the topology on $\mathcal{U}$ is exactly the quotient (final) topology induced by the encoders, allowing the latent overlaps to glue the components together without introducing spurious identifications or topological pathologies.
Let $\pi:X\to X/\!\sim$ be the quotient map. Since $E$ is constant on $\sim$-equivalence classes by definition, the universal property of quotient spaces yields a unique continuous map $\widetilde{E}:X/\!\sim\ \to\ \mathcal{U}$ such that
$E=\widetilde{E}\circ\pi.$
Bijectivity. By construction, $\mathcal{U}=\bigcup_{k=1}^{K}E^{[k]}(\mathcal{M}_{x}^{[k]})=E(X)$, so $\widetilde{E}$ is surjective. If $\widetilde{E}([x])=\widetilde{E}([x^{\prime}])$, then $E(x)=E(x^{\prime})$, hence $x\sim x^{\prime}$ and $x]=[x^{\prime}]$; thus $\widetilde{E}$ is injective.
$E$ is a quotient map. We show that the global encoder $E:X\to\mathcal{U}$ is a quotient map. Let $V\subseteq\mathcal{U}$. Since $X$ has the coproduct topology, $E^{-1}(V)$ is open in $X$ if and only if $(E^{[k]})^{-1}(V)$ is open in $\mathcal{M}_{x}^{[k]}$ for all $k$. By the coherence condition (ii), this holds if and only if $V$ is open in $\mathcal{U}$ (with the subspace topology). Therefore,
[…]
which is precisely the definition of $E$ being a quotient map.
$\widetilde{E}$ is a quotient map. Although quotientness does not in general pass to factors of a composition, we can verify directly that $\widetilde{E}$ is a quotient map. For any $V\subseteq\mathcal{U}$, using $E^{-1}(V)=\pi^{-1}(\widetilde{E}^{-1}(V))$ and the fact that $\pi$ is a quotient map, we have
[…]
Therefore $\widetilde{E}$ is a quotient map.
Conclusion. We have shown that $\widetilde{E}:X/\!\sim\ \to\ \mathcal{U}$ is a bijective quotient map. A standard result in topology states that a bijective quotient map is a homeomorphism. Therefore $\widetilde{E}$ is a homeomorphism. ∎
Remark 2 (Coherence as a quotient-map property).
Condition (ii) is equivalent to requiring that the global encoder $E:X\to\mathcal{U}$ is a quotient map (with $\mathcal{U}$ carrying the subspace topology from $\mathbb{R}^{d}$). Unlike an “open cover” assumption, (ii) allows the latent patches $\mathcal{M}_{z}^{[k]}$ to be closed or to overlap only along boundaries, which is common in applications.
A.2 Proof of Theorem 2 (Bi-Lipschitz Sufficiency)
Theorem 2. Let $(\mathcal{M}_{x},d_{X})$ and $(\mathcal{M}_{z},d_{Z})$ be metric spaces, each equipped with the topology induced by its metric. If the encoder $E:\mathcal{M}_{x}\to\mathcal{M}_{z}$ is $(c_{1},c_{2})$-bi-Lipschitz with $c_{1}>0$, then $E$ is a homeomorphism onto its image. Thus, constraining the Lipschitz constants satisfies Condition (i) of Theorem 1.